the length of a rectangle is 5m more than twice the width. The area is 42m(squared). Find length and width
Answer by checkley77(12844)

L=2W+5
LW=42
(2W+5)W=42
2W^2+5W-42=0
(2W-7)(W+6)=0
2W-7=0
2W=7
W=3.5 ANS.
L=2*3.5+5=7+5=12 ANS.
PROOF:
3.5*12=42
42=42